The equation that I am dealing with says
6/(z-5)-60/(z^2-25)=1
I have tried to solve this several times not and I come up with an answer that I am really not able to factor and solve. any help would be a blessing right now! thank you!

Answer by checkley79(3341)

6/(z-5)-60/(z^2-25)=1
6/(Z-5)-60/(Z+5)(Z-5)=1
[6(Z+5)-60]/(Z+5)(Z-5)=1
[6Z+30-60]/(Z+5)(Z-5)=1
[6Z-30]/(Z+5)(Z-5)=1 CROSS MULTIPLY.
(Z+5)(Z-5)=6Z-30
(Z+5)(Z-5)=6(Z-5) DIVIDE BY (Z-5).
Z+5=6
Z=6-5
Z=1 ANS..
PROOF:
6/(1-5)-60/(1^2-25)=1
6/-4-60/(1-25)=1
-1.5-60/-24=1
-1.5+2.5=1
1=1